6. Analyze the two economic alternatives described below and seleot the annual interest rate of 7% for both alternatives. All values are in$. best one. Use an Alternative Initial Cost Yearly Operating Expenses Annual Revenues Salvage Value Life (years) 22,000 10,000 2,000 3,000 6,000 10.000 2
Answer #1

We can do the analysis of best alternative based on present worth analysis. We calculate the present worth of alternative A. PW of A = -10,000 - 2000(P/A,i, n) + 3000(P/A, i, n) +2000(P/F, i,n) . Here I=7% and n=infinity for alternative A. Therefore PW of A = -10,000 -2000( P/A, 7%, infinity) + 3000(P/A, 7%, infinity) + 2000(P/F, 7%, infinity) = -10000 -2000/0.07 +3000/0.07 +0. As n=infinity salvage value at present period is negligible. So PW of A = -10,000 -2000/0.07+3000/0.07 = $4285.71 =$ 4286.

Similarly we can calculate PW of B is = -22000 +6000(P/A, 7%, 2) +10000(P/F, 7%, 2) = -10000 + 6000*1.808 +10000*0.8734 = $9582. (P/A, i, n) = {(1+0.07)^2 -0.07}/{0.07(1.07)^2}=1.808, (P/F,I,n) = (1.07)^-2=0.8734

So by comparing alternative A and B we can say present worth of A is less than PW of B. As B's PW is more ($9582) than A($4286) so we can say B is best alternative.